OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 5x – 2y – z = 16 x + y + 4z = 6 3x + 2y + 6z = 16 Choices: A)–7y – 21z = –14 –y – 6z = –2 B)7x + 7z = 28 x – 2z = 4 C)–3y + 19z = –14 5y + 18z = –2 D)21x – 7y = 70 33x – 10y = 112

OpenStudy (anonymous):

The answer is A. if you isolate x in eq2, you get x = (6 - y - 4z)... plug that into eq1 and you get... 5(6 - y - 4z) -2y -z = 16... 30 -5y -20z -2y -z = 16... -7y - 21z = 16-30 = -14... -7y - 21z = -14... Now plugging (6 - y - 4z) into eq3... 3(6 - y - 4z) +2y +6z = 16 18 - 3y -12z +2y +6z = 16 18 -1y -6z = 16 -y-6z = 16 - 18 = -2 -y-6z = -2

OpenStudy (kinggeorge):

A. All you have to do is R1 - 5xR2 and R3 - 3xR2. R1, R2, and R3 are equation 1, 2, and 3 respectively. You multiply R2 by 5 in the first because R1 has a 5x, and R2 only has x. The same reason applies for the second subtraction.

OpenStudy (anonymous):

ok thank u :)

OpenStudy (anonymous):

Hmmm. Is that the same as the substitution method?

OpenStudy (kinggeorge):

Good point, I didn't read the problem carefully enough. You can still stop once you reach -7y - 21z = -14 though.